# Effort: Efficient Orthogonal Modeling for Generalizable AI-Generated Image Detection

### v1

@@ 4 Experiment
F3Net [64] 22M 0.789 0.844 0.718 0.749 0.730 0.728 0.649 0.743 0.809 0.808 0.867 0.494 0.717 0.845 0.757 0.674 0.746
SPSL [47] 21M 0.799 0.871 0.724 0.770 0.723 0.702 0.794 0.769 0.747 0.748 0.885 0.514 0.666 0.812 0.643 0.665 0.710
SRM [53] 55M 0.840 0.885 0.695 0.728 0.722 0.702 0.794 0.767 0.749 0.704 0.779 0.704 0.659 0.772 0.793 0.694 0.732
CORE [58] 22M 0.809 0.882 0.721 0.720 0.765 0.724 0.710 0.762 0.871 0.843 0.959 0.679 0.774 0.958 0.855 0.724 0.833
RECCE [4] 48M 0.823 0.891 0.696 0.734 0.784 0.756 0.711 0.779 0.898 0.832 0.925 0.683 0.848 0.949 0.848 0.768 0.844
SLADD [8] 21M 0.837 0.904 0.772 0.756 0.800 0.690 0.683 0.777 0.878 0.882 0.954 0.765 0.825 0.943 0.879 0.794 0.865
SBI [72] 18M 0.886 0.827 0.717 0.848 0.899 0.703 0.866 0.821 0.724 0.891 0.952 0.750 0.594 0.803 0.712 0.701 0.766
UCF [91] 47M 0.837 0.867 0.742 0.770 0.808 0.774 0.697 0.785 0.831 0.827 0.950 0.731 0.862 0.937 0.809 0.647 0.824
IID [30] 66M 0.838 0.939 0.700 0.689 0.808 0.666 0.762 0.789 0.839 0.789 0.888 0.766 0.844 0.927 0.789 0.644 0.811
TALL [90] 87M 0.831 0.833 0.693 0.739 0.793 0.673 0.679 0.749 0.714 0.699 0.805 0.651 0.768 0.863 0.762 0.616 0.735
@@
Table 1: Benchmarking Results of Cross-dataset Evaluations (Protocol-1) and Cross-method Evaluations (Protocol-2). All detectors are trained on FF++_c23 [70] and evaluated on other fake data. The best results are highlighted in bold and the second is underlined. $\dagger$ indicates the results are obtained by using the model’s checkpoint provided by the authors, otherwise, the results are cited from [92, 94, 11].
@@
CNN-Spot [86] 100.0 93.47 84.50 99.54 89.49 98.15 89.02 73.75 59.47 98.24 98.40 73.72 70.62 71.00 70.54 80.65 84.91 82.07 70.59 83.58
Patchfor [5] 80.88 72.84 71.66 85.75 65.99 69.25 76.55 76.19 76.34 74.52 68.52 75.03 87.10 86.72 86.40 85.37 83.73 78.38 75.67 77.73
Co-occurence [57] 99.74 80.95 50.61 98.63 53.11 67.99 59.14 68.98 60.42 73.06 87.21 70.20 91.21 89.02 92.39 89.32 88.35 82.79 80.96 78.11
Freq-spec [98] 55.39 100.0 75.08 55.11 66.08 100.0 45.18 47.46 57.12 53.61 50.98 57.72 77.72 77.25 76.47 68.58 64.58 61.92 67.77 66.21
@@
CNN-Spot [86] 99.99 85.20 70.20 85.70 78.95 91.70 53.47 66.67 48.69 86.31 86.26 60.07 54.03 54.96 54.14 60.78 63.80 65.66 55.58 69.58
Patchfor [5] 75.03 68.97 68.47 79.16 64.23 63.94 75.54 75.14 75.28 72.33 55.30 67.41 76.50 76.10 75.77 74.81 73.28 68.52 67.91 71.24
Co-occurence [57] 97.70 63.15 53.75 92.50 51.10 54.70 57.10 63.06 55.85 65.65 65.80 60.50 70.70 70.55 71.00 70.25 69.60 69.90 67.55 66.86
Freq-spec 49.90 99.90 50.50 49.90 50.30 99.70 50.10 50.00 48.00 50.60 50.10 50.90 50.40 50.40 50.30 51.70 51.40 50.40 50.00 55.45
@@ 4.1 Deepfake Image Detection
Implementation Details. We utilize CLIP ViT-L/14 [65] as the default vision foundation model (VFM). We also investigate other VFMs in Tab. 5. We follow the pre-processing and training pipeline and use the codebases of DeepfakeBench [92]. Additionally, we sample 8 frames from each video for training and 32 frames for inference, following [72]. We use the fixed learning rate of 2e-4 for training our approach and employ the Adam [36] for optimization. We set the batch size to 32 for both training and testing. We also employ several widely used data augmentations, such as Gaussian Blur and Image Compression, following other existing works [93, 72, 11]. For the evaluation metric, we report the widely-used video-level Area Under the Curve (AUC) to compare our approach with other works, following [46, 90, 72]. Similar to these works, we compute the average model’s output probabilities of each video to obtain the video-level AUC.
@@
We adopt two widely used and standard protocols for evaluation: Protocol-1: cross-dataset evaluation and Protocol-2: cross-manipulation evaluation within the FF++ data domain. For Protocol-1, we conduct evaluations by training the models on FaceForensics++ (FF++) [70] and testing them on other seven deepfake detection datasets: Celeb-DF-v2 (CDF-v2) [45], DeepfakeDetection (DFD) [16], Deepfake Detection Challenge (DFDC) [15], the preview version of DFDC (DFDCP) [19], DeeperForensics (DFo) [32], WildDeepfake (WDF) [105], and FFIW [101]. Note that FF++ has three different compression versions and we adopt the c23 version for training all methods in our experiments, following most existing works [93]. For Protocol-2, we evaluate the models on the latest deepfake dataset DF40 [95], which contains the forgery data generated within the FF++ domain, ensuring the evaluation fake methods different but the data domain unchanged. In this manner, we can assess the model’s generalization when facing unseen fake methods, ignoring the impact of the data domain gap.
@@
To provide a comprehensive benchmark for comparison, we introduce 13 competitive detectors, including several classical detection methods such as F3Net [64] (ECCV’20), SPSL [47] (CVPR’20), SRM [47] (CVPR’21), CORE [58] (CVPRW’22), RECCE [4] (CVPR’22), and SBI [72] (CVPR’22), and also several latest SOTA methods (after 2023), such as UCF [91] (ICCV’23), IID [30] (CVPR’23), TALL [90] (ICCV’23), LSDA [94] (CVPR’24), ProDet [11] (NeurIPS’24), and CDFA [46] (ECCV’24). All detectors are trained on FF++ (c23) and tested on other fake data. Results in Tab. 1 demonstrate two notable advantages of our approach. (1) generalizability: we see that our approach consistently and largely outperforms other models across basically all tested scenarios, validating that our method is generalizable for detecting unseen fake data, even for the latest face-swapping techniques such as BleFace [73]. (2) efficiency: it is worth noting that our method only needs 0.19M parameters for training to achieve superior generalization. As we can see most latest SOTA detectors such as LSDA and ProDet all use about 100M parameters for training, while we are about 1,000$\times$ smaller.
@@ 4.2 Synthetic Image Detection
We follow existing works [85, 61, 48] for benchmarking and report both average precision (AP) and classification accuracy (Acc). For Acc, we set the classification threshold for each dataset to $0.5$ to ensure a fair comparison.
Table 4: Ablation studies regarding the tunable $n-r$ values in SVD. All models are trained on FF++ (c23) and tested on CDF-v2 and SimSwap. “FFT” indicates the full fine-tuning. “Linear-Prob” indicates fine-tuning the FC layer only.
Archs. #Params $n-r$ CDF-v2 SimSwap Avg.
Baseline (Linear-Prob) 0.002M 0.765 0.769 0.767
Baseline (FFT) 307M 0.857 0.860 0.859
Ours Variant-1 12M 64 0.946 0.935 0.941
Variant-2 3M 16 0.948 0.930 0.939
Variant-3 0.75M 4 0.955 0.921 0.938
Variant-4 0.19M 1 0.956 0.926 0.941
@@
UniversalFakeDetect Dateset.
We adhere to the protocol outlined in [85, 61] and utilize ProGAN’s real and fake images as our training dataset, which includes 20 subsets of generated images. The evaluation set contains 19 subsets derived from different kinds of generative models, including ProGAN [33], CycleGAN [102], BigGAN [2], StyleGAN [34], GauGAN [62], StarGAN [13], DeepFakes [69], SITD [7], SAN [14], CRN [10], IMLE [43], Guided (guided diffusion model)[17], LDM (latent diffusion model) [67], Glide [60], and DALLE [66].
Implementation Details.
Similar to the setting of deepfake image detection, we adopt pre-trained CLIP ViT-L/14 as the backbone and use the Adam optimizer [36] with a fixed learning rate of 2e-4. The batch size is set to 48. Other settings and details are the same with [61].
@@
The AP and Acc results are presented in Tab. 2 and Tab. 3, respectively. Our method attains impressive detection results, achieving 95.19% mAcc and 99.41% mAP across the 19 test subsets. One similar approach to ours is UniFD, which also preserves the original semantic knowledge of CLIP and fine-tunes only the FC layer for discrimination. In contrast to UniFD which performs discrimination in the semantic space, our approach utilizes SVD to create an orthogonal low-ranked subspace for learning forgeries while preserving the essential high-ranked semantic space, achieving the discrimination by leveraging both semantic and forgery subspaces, achieving a significant improvement of 9.27% in mAcc and 13.81% in mAP over UniFD. Besides, when compared to the SoTA method, FatFormer, we achieve 4.33% mAcc improvement without relying on the extra text encoder of CLIP. This further demonstrates the superiority of our approach.
Table 6: Ablation studies regarding the singular value constraint ($\mathcal{L}_{\rm ksv}$) and orthogonal constraint ($\mathcal{L}_{\rm orth}$). All models are trained on FF++ (c23) and tested on other datasets.
Ours CDF-v2 DFDC SimSwap FSGAN Avg.
SVD $\mathcal{L}_{\rm ksv}$ $\mathcal{L}_{\rm orth}$
$\times$ $\times$ $\times$ 0.857 0.758 0.860 0.939 0.854
$\times$ $\times$ 0.940 0.829 0.910 0.955 0.909
$\times$ 0.944 0.841 0.927 0.953 0.916
$\times$ 0.945 0.847 0.914 0.950 0.914
0.956 0.843 0.926 0.957 0.921
@@ 4.3 Ablation Study and Analysis
How many singular values should be tuned when applying SVD?
When applying SVD, we create the semantic (frozen) and forgery (tuning) subspaces. But a crucial question arises: how many singular values should be set aside for freezing and how many for tuning? To figure out the optimal empirical values for this, we carry out an ablation study. The purpose of this study is to evaluate the effectiveness of different values of $n-r$. Here, $n$ represents the full-ranked dimension of the original weights, while $r$ is the low-ranked dimension that we intend to tune We examine four different variants with specific values for $n-r$, namely 64, 16, 6, and 1. Additionally, we include the baseline case, which is the full-finetune CLIP, for the sake of comparison. Results in Tab. 4 suggest that our approach largely outperforms the baseline. Moreover, the performance levels of the four variants are quite similar. This implies that none of these variants holds a distinct advantage when it comes to generalization. However, among them, variant-4 has the fewest trainable parameters. Consequently, we select this particular setting as the default option in our other experiments.
Figure 5: Evidence for verifying the illustration of Fig. 2, showing the latent distributions of vanilla CNN (e.g., Res50 [86]) and ours.
Figure 6: Evidence for validating the discrimination dimension from the logits space. Please refer to the text for details.
@@
By default, we choose CLIP as the vision foundation model (VFM) in our experiments. To validate the generality and versatility of our approach, we conduct an ablation study to apply our method to other VFMs, including BEIT-v2 [63] and SigLIP [96]. Results in Tab. 5 show that our approach can be seamlessly applied to other existing VFMs for improving the model’s generalization performance.
Figure 7: Evidence for validating the discrimination dimension from the feature space. We use PCA to analyze the rank of the feature space. Please refer to the text for details.
@@
How to verify the naively trained models use only forgery patterns for discrimination?
In our motivation (RQ-2), we argue that the naively trained models rely solely on forgery patterns for discrimination, as the models tend to overfit the seen fake patterns during training (see Fig. 1, Fig. 3, Fig. 5). To verify that the baseline models (a vanilla CNN, i.e., Xception) indeed leverages only the forgery patterns for discrimination, we conduct a new visualization in Fig. 6, providing reasonable evidence to see the model’s discrimination behavior. Specifically, we visualize the model’s output, i.e., logits, for both real and fake classes, and a dotted line for the boundary (to those samples with real logits equal to fake logits). As shown in Fig. 6, we see that all prediction samples of Xception lie in a linear line (similar to $y=-x+b$), indicating that the Xception relies on one dimension for discriminating real and fake. In addition, the CLIP contains rich semantic knowledge, but fully fine-tuning CLIP can distort the well-learned semantic knowledge, making it near to a line ($y=-x+b$). Finally, our approach maximizes the retention of semantic knowledge. We achieve this by learning forgeries within a new orthogonal subspace relative to the semantic subspaces, achieving superior generalization than the other two models.
Low-ranked feature space causes the model’s overfitting.
We use Principal Component Analysis (PCA) to visualize the feature space with the two most principal components (the top row) and demonstrate the explained variance ratio of different principal components (the below row), which is similar to “how much of the total information of the data is captured by each principal component”. Results in Fig. 7 show that the feature space of the Xception is highly low-ranked, with only two principal components to capture all information, resulting in limited generalization.

### v2

@@ 4 Experiment
Table 1: Benchmarking Results of Cross-dataset Evaluations (Protocol-1) and Cross-method Evaluations (Protocol-2). All detectors are trained on FF++_c23 (Rossler et al., 2019) and evaluated on other fake data. $\dagger$ indicates the results are obtained by using the model’s checkpoint provided by the authors, otherwise, the results are cited from (Yan et al., 2023b, 2024a; Cheng et al., 2024).
@@
F3Net (Qian et al., 2020) 22M 0.789 0.844 0.718 0.749 0.730 0.728 0.649 0.743 0.809 0.808 0.867 0.494 0.717 0.845 0.757 0.674 0.746
SPSL (Liu et al., 2021a) 21M 0.799 0.871 0.724 0.770 0.723 0.702 0.794 0.769 0.747 0.748 0.885 0.514 0.666 0.812 0.643 0.665 0.710
SRM (Luo et al., 2021) 55M 0.840 0.885 0.695 0.728 0.722 0.702 0.794 0.767 0.749 0.704 0.779 0.704 0.659 0.772 0.793 0.694 0.732
CORE (Ni et al., 2022) 22M 0.809 0.882 0.721 0.720 0.765 0.724 0.710 0.762 0.871 0.843 0.959 0.679 0.774 0.958 0.855 0.724 0.833
RECCE (Cao et al., 2022) 48M 0.823 0.891 0.696 0.734 0.784 0.756 0.711 0.779 0.898 0.832 0.925 0.683 0.848 0.949 0.848 0.768 0.844
SLADD (Chen et al., 2022a) 21M 0.837 0.904 0.772 0.756 0.800 0.690 0.683 0.777 0.878 0.882 0.954 0.765 0.825 0.943 0.879 0.794 0.865
SBI (Shiohara & Yamasaki, 2022) 18M 0.886 0.827 0.717 0.848 0.899 0.703 0.866 0.821 0.724 0.891 0.952 0.750 0.594 0.803 0.712 0.701 0.766
UCF (Yan et al., 2023a) 47M 0.837 0.867 0.742 0.770 0.808 0.774 0.697 0.785 0.831 0.827 0.950 0.731 0.862 0.937 0.809 0.647 0.824
IID (Huang et al., 2023) 66M 0.838 0.939 0.700 0.689 0.808 0.666 0.762 0.789 0.839 0.789 0.888 0.766 0.844 0.927 0.789 0.644 0.811
@@
CNN-Spot (Wang et al., 2020b) 100.0 93.47 84.50 99.54 89.49 98.15 89.02 73.75 59.47 98.24 98.40 73.72 70.62 71.00 70.54 80.65 84.91 82.07 70.59 83.58
Patchfor (Chai et al., 2020) 80.88 72.84 71.66 85.75 65.99 69.25 76.55 76.19 76.34 74.52 68.52 75.03 87.10 86.72 86.40 85.37 83.73 78.38 75.67 77.73
Co-occurence (Nataraj et al., 2019) 99.74 80.95 50.61 98.63 53.11 67.99 59.14 68.98 60.42 73.06 87.21 70.20 91.21 89.02 92.39 89.32 88.35 82.79 80.96 78.11
Freq-spec (Zhang et al., 2019) 55.39 100.0 75.08 55.11 66.08 100.0 45.18 47.46 57.12 53.61 50.98 57.72 77.72 77.25 76.47 68.58 64.58 61.92 67.77 66.21
@@
CNN-Spot (Wang et al., 2020b) 99.99 85.20 70.20 85.70 78.95 91.70 53.47 66.67 48.69 86.31 86.26 60.07 54.03 54.96 54.14 60.78 63.80 65.66 55.58 69.58
Patchfor (Chai et al., 2020) 75.03 68.97 68.47 79.16 64.23 63.94 75.54 75.14 75.28 72.33 55.30 67.41 76.50 76.10 75.77 74.81 73.28 68.52 67.91 71.24
Co-occurence (Nataraj et al., 2019) 97.70 63.15 53.75 92.50 51.10 54.70 57.10 63.06 55.85 65.65 65.80 60.50 70.70 70.55 71.00 70.25 69.60 69.90 67.55 66.86
Freq-spec (Zhang et al., 2019) 49.90 99.90 50.50 49.90 50.30 99.70 50.10 50.00 48.00 50.60 50.10 50.90 50.40 50.40 50.30 51.70 51.40 50.40 50.00 55.45
@@ 4.1 Deepfake Image Detection
Implementation Details. We utilize CLIP ViT-L/14 (Radford et al., 2021) as the default vision foundation model (VFM). We also investigate other VFMs in Tab. 5. We follow the pre-processing and training pipeline and use the codebases of DeepfakeBench (Yan et al., 2023b). Additionally, we sample 8 frames from each video for training and 32 frames for inference, following (Shiohara & Yamasaki, 2022). We use the fixed learning rate of 2e-4 for training our approach and employ the Adam (Kingma & Ba, 2014) for optimization. We set the batch size to 32 for both training and testing. We also employ several widely used data augmentations, such as Gaussian Blur and Image Compression, following other existing works (Yan et al., 2024a; Shiohara & Yamasaki, 2022; Cheng et al., 2024). For the evaluation metric, we report the widely-used video-level Area Under the Curve (AUC) to compare our approach with other works, following (Lin et al., 2024; Xu et al., 2023; Shiohara & Yamasaki, 2022). Similar to these works, we compute the average model’s output probabilities of each video to obtain the video-level AUC.
@@
We adopt two widely used and standard protocols for evaluation: Protocol-1: cross-dataset evaluation and Protocol-2: cross-manipulation evaluation within the FF++ data domain. For Protocol-1, we conduct evaluations by training the models on FaceForensics++ (FF++) (Rossler et al., 2019) and testing them on other seven deepfake detection datasets: Celeb-DF-v2 (CDF-v2) (Li et al., 2020b), DeepfakeDetection (DFD) (DFD., 2020), Deepfake Detection Challenge (DFDC) (detection challenge., 2020), the preview version of DFDC (DFDCP) (Dolhansky et al., 2019), DeeperForensics (DFo) (Jiang et al., 2020), WildDeepfake (WDF) (Zi et al., 2020), and FFIW (Zhou et al., 2021). Note that FF++ has three different compression versions and we adopt the c23 version for training all methods in our experiments, following most existing works (Yan et al., 2024a). For Protocol-2, we evaluate the models on the latest deepfake dataset DF40 (Yan et al., 2024b), which contains the forgery data generated within the FF++ domain, ensuring the fake methods different while the data domains remain unchanged.
@@
To provide a comprehensive benchmark for comparison, we introduce 13 competitive detectors. All detectors are trained on FF++ (c23) and tested on other fake data. Results in Tab. 1 demonstrate two notable advantages of our approach. (1) generalizability: we see that our approach consistently and largely outperforms other models across basically all tested scenarios, validating that our method is generalizable for detecting unseen fake data, even for the latest face-swapping techniques such as BleFace (Shiohara et al., 2023). (2) efficiency: it is worth noting that our method only needs 0.19M parameters for training to achieve superior generalization. As we can see most latest SOTA detectors such as LSDA and ProDet all use about 100M parameters for training, while we are about 1,000$\times$ smaller.
@@ 4.2 Synthetic Image Detection
We follow existing works (Wang et al., 2020a; Ojha et al., 2023; Liu et al., 2024) for benchmarking and report both average precision (AP) and classification accuracy (Acc). For Acc, we set the classification threshold for each dataset to $0.5$ to ensure a fair comparison.
UniversalFakeDetect Dateset.
We adhere to the protocol outlined in (Wang et al., 2020a; Ojha et al., 2023) and utilize ProGAN’s real and fake images as our training dataset, which includes 20 subsets of generated images. The evaluation set contains 19 subsets derived from different kinds of generative models, including ProGAN (Karras et al., 2018), CycleGAN (Zhu et al., 2017), BigGAN (Brock et al., 2018a), StyleGAN (Karras et al., 2019), GauGAN (Park et al., 2019), StarGAN (Choi et al., 2018), DeepFakes (Rössler et al., 2019), SITD (Chen et al., 2018), SAN (Dai et al., 2019), CRN (Chen & Koltun, 2017), IMLE (Li et al., 2019), Guided (guided diffusion model)(Dhariwal & Nichol, 2021), LDM (latent diffusion model) (Rombach et al., 2022a), Glide (Nichol et al., 2022), and DALLE (Ramesh et al., 2021).
Implementation Details.
Similar to the setting of deepfake image detection, we adopt pre-trained CLIP ViT-L/14 as the backbone and use the Adam optimizer (Kingma & Ba, 2014) with a fixed learning rate of 2e-4. The batch size is set to 48. Other settings and details are the same with (Ojha et al., 2023).
Table 4: Ablation studies regarding the proposed SVD, singular value constraint ($\mathcal{L}_{\rm ksv}$), and orthogonal constraint ($\mathcal{L}_{\rm orth}$). All models are trained on FF++ (c23) and tested on other datasets.
Ours CDF-v2 SimSwap Avg.
SVD $\mathcal{L}_{\rm ksv}$ $\mathcal{L}_{\rm orth}$
$\times$ $\times$ $\times$ 0.857 0.860 0.859
$\times$ $\times$ 0.940 0.910 0.925
$\times$ 0.944 0.927 0.936
$\times$ 0.945 0.914 0.930
0.956 0.926 0.941
@@
The AP and Acc results are presented in Tab. 2 and Tab. 3, respectively. Our method attains impressive detection results, achieving 95.19% mAcc and 99.41% mAP across the 19 test subsets. One similar approach to ours is UniFD, which also preserves the original pre-trained knowledge of CLIP and fine-tunes only the FC layer for discrimination. In contrast to UniFD which directly performs discrimination in the pre-trained knowledge space, our approach utilizes SVD to create an orthogonal low-ranked subspace for learning forgeries while preserving the essential high-ranked representational space, achieving the discrimination by leveraging both representational and forgery subspaces, achieving a significant improvement of 9.27% in mAcc and 13.81% in mAP over UniFD. Besides, when compared to the SOTA method, FatFormer, we achieve 4.33% mAcc improvement without relying on the extra text encoder of CLIP. This further demonstrates the superiority of our approach.
Table 6: Ablation studies on synthetic image detection regarding the tunable $n-r$ values in SVD (Ours) and $r$ values in LoRA. All models are trained on ProGAN’s images and tested on 19 different generative models’ images. “FFT” indicates the full fine-tuning. “Linear-Prob” indicates fine-tuning FC layer only, where we reproduce the results from UniFD (Ojha et al., 2023).
Archs. $n-r$ $r$ mAcc
UniFD (Linear-Prob) 81.02
Baseline (FFT) 86.22
LoRA Variant-1 256 91.42
Variant-2 64 91.06
Variant-3 16 91.89
Variant-4 4 93.53
Variant-5 1 93.03
Ours Variant-1 256 92.13
Variant-2 64 93.68
Variant-3 16 94.45
Variant-4 4 94.37
Variant-5 1 95.19
Table 7: Cross-dataset generalization evaluations with existing adapter-based deepfake detectors. The results are cited from their original papers. The metric is frame-level AUC.
Methods CDF-v2 DFD DFDC Avg.
LoRA (Kong et al., 2023) 0.838 0.834 0.717 0.796
MoE-LoRA (Kong et al., 2024) 0.867 0.904
Dual-Adapter (Shao et al., 2023) 0.717 0.727
Ours 0.901 0.923 0.798 0.874
@@ 4.3 Ablation Study and Analysis
Incremental improvement of the proposed designs.
Ablation studies in Tab. 4 demonstrate incremental gains from the proposed SVD method and loss constraints: the baseline (no modules) achieves an average AUC of 0.854, while adding SVD alone boosts performance to 0.909 (+6.5%), underscoring its efficacy in isolating forgery artifacts via orthogonal feature decomposition. Further integrating the singular value constraint ($\mathcal{L}_{\rm ksv}$) and orthogonal constraint ($\mathcal{L}_{\rm orth}$) refines performance, yielding 0.916 and 0.914 AUC, respectively, with their combined synergy (SVD + $\mathcal{L}_{\rm ksv}$ + $\mathcal{L}_{\rm orth}$) achieving peak performance (0.921 AUC). These results highlight the significant contribution of SVD as the primary driver of performance gains while illustrating the complementary benefits of $\mathcal{L}_{\rm ksv}$ and $\mathcal{L}_{\rm orth}$ in further refining the generalization performance.
@@
By default, we choose CLIP as the vision foundation model (VFM) in our experiments. To validate the generality and versatility of our approach, we conduct an ablation study to apply our method to other VFMs, including BEIT-v2 (Peng et al., 2022) and SigLIP (Zhai et al., 2023). Results in Tab. 5 show that our approach can be seamlessly applied to other VFMs to improve the model’s generalization performance.
Figure 5: PCA quantifying feature space information retention, measured by the minimum number of components required to explain $\geq 90\%$ variance. Our SVD-based fine-tuning preserves 159/160 principal components (retaining 99.3% original variance), while LoRA and FFT exhibit notable information degradation, with 14.4% and 64.4% variance loss, respectively.
Comparison with Existing Adapter-Based Detectors.
Parameter-efficient fine-tuning (PEFT) has become a popular technique for adapting pre-trained large models to downstream tasks (Ding et al., 2023). Low-ranked adaptation (LoRA) (Hu et al., 2021) is a widely used approach for PEFT. Previous works (Kong et al., 2023; Liu et al., 2024) employing LoRA in the VFMs have achieved good empirical generalization results for detection. Additionally, (Kong et al., 2024) and (Shao et al., 2023) introduce MoE-LoRA techniques and dual adapters into the deepfake detection fields. However, the existing adapter-based methods do not explicitly ensure this orthogonality, still having the potential to distort the pre-existing pre-trained knowledge and result in unexpected generalization results. In contrast, our method explicitly constructs two orthognoal subspaces for pre-trained knowledge and forgery using SVD, ensuring the pre-existing pre-trained knowledge will not be distorted, thereby achieving better generalization performance. To verify this, we provide several empirical results in Tab. 7 and Tab. 6. From these results, we can see that our proposed SVD-based method achieves clearly higher generalization results than adapter-based methods in both deepfake detection and synthetic image detection fields, as our approach explicitly preserves the pre-trained knowledge while learning the forgery patterns. We also use PCA to compute the rank of the feature space, similar to Fig. 3. Results in Fig. 5 highlight the superiority of our method, where full-parameters fine-tuning and LoRA can lead to a notable reduction (-103 and -23, respectively) while our method best retains the pre-trained knowledge, maintaining a higher-ranked feature space for better generalization.
Figure 6: Evidence for validating the discrimination dimension from the logits space. The first row shows that the vanilla CNN (i.e., Xception) overfits to the seen fakes and relies only on the forgery pattern for discrimination. The second row shows that fully fine-tuning the VFM (i.e., CLIP) involves pre-trained knowledge but also distorting part of it during learning forgery patterns. The third row shows that our method uses SVD to achieve the orthogonality, thereby best retaining the pre-trained knowledge.
@@
The naively trained models leverage very limited forgery patterns for discrimination.
In our motivation, we argue that conventional training paradigms cause models to over-rely on very limited forgery patterns for discrimination, thereby causing the highly low-ranked and constrained feature space, limiting their expressivity and generalization (Fig. 2 and Fig. 3). To further validate this argument, we analyze the discrimination behavior of baseline models (e.g., Xception and CLIP) through decision boundary visualization using the model’s output logits of real and fake classes (Fig. 6). The linear alignment of predictions along $y=-x+b$ reveals that Xception collapses real/fake discrimination into a single discriminative dimension, confirming its exclusive dependence on forgery cues. Notably, while CLIP’s original higher-ranked representations initially preserve informative structures, full fine-tuning catastrophically degrades this structure, forcing decisions into a similarly collapsed subspace ($y=-x+b$). Our method addresses this concern by learning forgery-related features in a novel orthogonal complement subspace relative to CLIP’s original representational embedding space, thereby achieving optimal generalization performance.